# Low-Altitude Wireless Networks: A Survey

### v1

@@ 6.1 LAA Structuring
4) Blocky Airspace: Block-shaped airspace refers to the division of LAA into horizontal blocks to align with urban structures or specific functional requirements. These blocks differ in infrastructure technology and service provider configurations. Compared to corridor and layered airspace, block-shaped airspace offers greater flexibility, with fewer operational restrictions, allowing more efficient use of airspace. However, block-shaped airspace does not inherently provide aircraft separation, incurring significant challenges to effective airspace management and conflict avoidance. To achieve this, the SESAR CORUS project [158] introduced X, Y, and Z block-shaped airspaces by jointly considering environmental/social factors and the offered service priority, as shown in Fig. 6(d). X airspace lies in the lowest service level with aircraft with low task risk, e.g., rural and sportful areas. By contrast, Y airspace is a higher level with strategic conflict avoidance services like those in urban areas. Z airspace is on the top, which supports larger traffic flows and mitigates extensive risk, e.g., the city center and near airports.
@@
5) Free Airspace: Different from the above structured LAA, free airspace is a non-structured framework that does not have fixed flight paths or divisions in either the horizontal or vertical direction. In the early 21st century, the national aerospace laboratory (NLR) of the Netherlands proposed free flight in the civil aviation domain [159]. The core objective is to enable aircraft to autonomously separate and avoid conflicts through the system, allowing enhanced freedom to select preferred flight paths. Based on this design, the Metropolis project [153] proposed free LAA illustrated in Fig. 6(e). This mechanism only specifies the allowed flight altitude and takes responsibility for managing airspace capacity and setting airspace boundaries. The collision avoidance and aircraft separation are fully addressed by the drones themselves. Due to its unstructured nature, the capacity of free LAA is significantly enhanced. More importantly, the drones only face minimal flight restrictions, and thus significantly shorten their flight distances to the destinations. However, the successful application of free airspace requires a well-developed collision avoidance mechanism. As the missions become complicated, ensuring real-time collision avoidance in dense LAWN deployments remains an open issue.
6.2 Low-Altitude Air Traffic Management
@@
Apart from effective LAA structuring, the management of LAA has become a critical issue with the rise of urban air mobility. To address the challenges presented by growing air traffic densities, various ATM methods have been proposed. One prominent approach is the development of macroscopic fundamental diagrams (MFD), traditionally used in ground traffic management [160]. Specifically, MFDs characterize the relationship between traffic flow, density, and speed at the network level. Empirical studies have demonstrated that, similar to road traffic, the flow of air traffic decreases as density increases, as illustrated in Fig. 7. It can be observed that as traffic density increases, the traffic flow initially rises, reaching a peak before dropping off due to congestion. The transition from free flow to congestion occurs when density exceeds the optimal level, highlighting the LAA’s capacity limit. However, it is noteworthy that, even at high densities, air traffic flow does not always decrease by leveraging active management strategies such as departure/landing sequencing and air traffic control measures. The work of [161] presented a throughput-based capacity metric to evaluate the impact of increasing air traffic on LAA. The proposed method simulated traffic and analyzed the effects of different conflict resolution methods, where the findings suggest that throughput decreases as traffic density rises, even while maintaining safety, offering a new way to assess conflict detection strategies. The authors of [162] addressed the challenges faced by the national airspace system in optimizing traffic flow, in which various strategies for managing air traffic are reviewed. The authors emphasized the role of collaborative decision-making (CDM) and automation to improve capacity and reduce congestion in high-density environments. Furthermore, the study of [10] introduced a new approach to air traffic flow optimization using convex formulations of partial differential equations (PDEs). The authors modeled air traffic flow in an Eulerian network and demonstrated how to convert these models into linear and quadratic optimization problems, providing globally optimal solutions for large-scale air traffic systems.
@@
A machine learning-based decision support system (DSS) for ATM was further developed by integrating explainable AI (XAI) in [163]. By employing learning-based algorithms to predict risks, this approach can assist controllers in managing air traffic effectively. The incorporation of XAI ensures that the reasoning behind AI decisions is transparent and understandable to human operators, thereby increasing trust in automated systems. Furthermore, the research explored the system’s application in both controlled and uncontrolled airspace, illustrating its potential to enhance safety and operational efficiency in low-altitude ATM. To achieve efficient ATM in real-world environments, significant reforms are being made in the strategic planning and development of ATM systems under the single European sky (SES) initiative [164], motivated by the need to establish unified strategic objectives to improve LAWNs’ capacity, safety, and environmental sustainability. A comprehensive review of the ATM planning framework in Europe has revealed critical gaps, particularly regarding the alignment of performance targets with local operational needs. The findings suggest that current strategic goals for capacity are overly broad and fail to address the specific requirements of individual regions and airspaces. To sum up, although the low-altitude ATM development is progressing, it remains in the early stages to support practical implementation.
@@ 7.1 Digital Twin-Enabled LAWNs
7.2 Lightweight Large Language Models-Enhanced LAWNs
Deploying large language models (LLMs) within LAWNs can facilitate intelligent decision-making, autonomous mission planning, and semantic-aware resource allocation directly on UAVs or edge nodes. This paradigm reduces the dependence on centralized cloud computing, alleviates latency, and improves responsiveness, which is critical for mission-critical applications, such as disaster response, traffic monitoring, and logistics. Nevertheless, the deployment of LLMs at the network edge is impeded by constraints on computational resources, energy consumption, and model compression. To address these limitations, the development of lightweight, energy-efficient LLM architectures and distributed inference frameworks tailored for mobile aerial platforms remains a key open challenge for next-generation intelligent LAWNs.
@@ 7.3 Standardizations for LAWN Deployments
Despite the significant potential of LAWNs in enabling low-altitude economy applications, their large-scale adoption is hindered by the lack of standardized frameworks and regulatory guidelines. Key challenges include defining airspace usage protocols, ensuring interoperability among heterogeneous UAV platforms, allocating spectrum for aerial communications, and ensuring compliance with safety and privacy regulations. Standardization efforts must strike a balance between promoting innovation and ensuring regulatory compliance, enabling the reliable, scalable, and secure deployment of LAWNs. Furthermore, the global harmonization of technical, operational, and legal requirements across various regions remains a critical issue that must be addressed to fully unlock the commercial potential of LAWNs.
@@ 8 Conclusion
This paper provided a comprehensive survey of LAWNs, focusing on their functional designs and LAA management. We began by comparing LAWNs with traditional aerial communication systems, highlighting their ability to integrate communication, sensing, computation, and control. We then established the foundational principles for LAWN architectures and channel modeling, followed by a review of recent advancements in functional designs and performance metrics. Finally, we summarized methodologies for LAA structuring and air traffic management, essential for ensuring scalable and safe drone operations.

### v2

@@ 7.1 LAA Structuring
4) Blocky Airspace: Block-shaped airspace refers to the division of LAA into horizontal blocks to align with urban structures or specific functional requirements. These blocks differ in infrastructure technology and service provider configurations. Compared to corridor and layered airspace, block-shaped airspace offers greater flexibility, with fewer operational restrictions, allowing more efficient use of airspace. However, block-shaped airspace does not inherently provide aircraft separation, incurring significant challenges to effective airspace management and conflict avoidance. To achieve this, the SESAR CORUS project [corus2019] introduced X, Y, and Z block-shaped airspaces by jointly considering environmental/social factors and the offered service priority, as shown in Fig. 8(d). X airspace lies in the lowest service level with aircraft with low task risk, e.g., rural and sportful areas. By contrast, Y airspace is a higher level with strategic conflict avoidance services like those in urban areas. Z airspace is on the top, which supports larger traffic flows and mitigates extensive risk, e.g., the city center and near airports.
Table 5: Comparison of Representative Airspace Types
Airspace Type Structure Capacity Safety Advantages & Disadvantages
Pipeline Airspace Virtual 3D zones Low High High control, low management cost; Limited flexibility, prone to congestion
Corridor Airspace Fixed parallel paths Moderate Moderate Fewer flight restrictions; Requires extensive planning and management at intersections
Layered Airspace Vertical layers Relatively high Moderate Reduces conflict, simple control; Reduces airspace efficiency, less freedom of use
Blocky Airspace Horizontal blocks High Relatively low Accommodate urban development; Complex management, no inherent separation of aircraft
Free Airspace Fully flexible Very high Low Minimal restrictions, high capacity; Requires advanced collision avoidance, difficult to manage in dense areas
@@
5) Free Airspace: Different from the above structured LAA, free airspace is a non-structured framework that does not have fixed flight paths or divisions in either the horizontal or vertical direction. In the early 21st century, the National Aerospace Laboratory (NLR) of the Netherlands proposed free flight in the civil aviation domain [hoekstra2001]. The core objective is to enable aircraft to autonomously separate and avoid conflicts through the system, allowing enhanced freedom to select preferred flight paths. Based on this design, the Metropolis project [metropolis2014] proposed free LAA illustrated in Fig. 8(e). This mechanism only specifies the allowed flight altitude and takes responsibility for managing airspace capacity and setting airspace boundaries. The collision avoidance and aircraft separation are fully addressed by the drones themselves. Due to its unstructured nature, the capacity of free LAA is significantly enhanced. More importantly, the drones only face minimal flight restrictions, and thus significantly shorten their flight distances to the destinations. However, the successful application of free airspace requires a well-developed collision avoidance mechanism. As the missions become complicated, ensuring real-time collision avoidance in dense LAWN deployments remains an open issue. For ease of exposition, we summarize the main features of the aforementioned airspace types in TABLE 5, which provides a concise comparison by focusing on their structure, capacity, safety considerations, and a balanced overview of their advantages and disadvantages.
7.2 Low-Altitude ATM
@@
Apart from effective LAA structuring, the management of LAA has become a critical issue with the rise of UAM. To address the challenges associated with increasing traffic densities, macroscopic fundamental diagrams (MFDs), originally developed for ground traffic management, have been adapted for LAA [Rastgoftar2024]. An MFD characterizes the aggregate relationship among traffic flow, density, and speed at the network level and thereby provides a compact description of the operational state of the airspace. As shown in Fig. 9, the simulation data and the fitted curve jointly reveal a characteristic three-regime behavior. In the free-flow regime, at low densities, the normalized traffic flow increases almost linearly with density. As density approaches a critical value, the flow reaches a maximum, which indicates the practical capacity of the airspace. Beyond this critical density, the system enters a congested regime, where additional aircraft lead to a reduction in flow and the emergence of unstable operating conditions. This relationship demonstrates that MFDs can capture capacity limits and congestion onset in LAA, and thus provide a useful tool for designing density thresholds, scheduling policies, and flow control strategies for LAWNs. The work of [Bulusu2018] presented a throughput-based capacity metric to evaluate the impact of increasing air traffic on LAA. The proposed method simulated traffic and analyzed the effects of different conflict resolution methods, where the findings suggest that throughput decreases as traffic density rises, even while maintaining safety, offering a new way to assess conflict detection strategies. The authors of [Li2022] addressed the challenges faced by the national airspace system in optimizing traffic flow, in which various strategies for managing air traffic are reviewed. The authors emphasized the role of collaborative decision-making (CDM) and automation to improve capacity and reduce congestion in high-density environments. Furthermore, the study of [Work2009] introduced a new approach to air traffic flow optimization using convex formulations of partial differential equations (PDEs). The authors modeled air traffic flow in an Eulerian network and demonstrated how to convert these models into linear and quadratic optimization problems, providing globally optimal solutions for large-scale air traffic systems.
@@
A machine learning-based decision support system for ATM was further developed by integrating explainable AI (XAI) in [aerospace8080224]. By employing learning-based algorithms to predict risks, this approach can assist controllers in managing air traffic effectively. The incorporation of XAI ensures that the reasoning behind AI decisions is transparent and understandable to human operators, thereby increasing trust in automated systems. Furthermore, the research explored the system’s application in both controlled and uncontrolled airspace, illustrating its potential to enhance safety and operational efficiency in low-altitude ATM. To achieve efficient ATM in real-world environments, significant reforms are being made in the strategic planning and development of ATM systems under the single European sky (SES) initiative [REZO20235], motivated by the need to establish unified strategic objectives to improve LAWNs’ capacity, safety, and environmental sustainability. A comprehensive review of the ATM planning framework in Europe has revealed critical gaps, particularly regarding the alignment of performance targets with local operational needs. The findings suggest that current strategic goals for capacity are overly broad and fail to address the specific requirements of individual regions and airspaces. To sum up, although the low-altitude ATM development is progressing, it remains in the early stages to support practical implementation.
Moreover, the deployment of LAWNs necessitates a strategic alignment with both international frameworks and domestic regulations to ensure safety, interoperability, and compliance. A comprehensive review of representative standards relevant to LAWNs is provided to demonstrate how such systems can seamlessly integrate into global and national UAV ecosystems while adhering to regulatory requirements. The ICAO has developed the ATM framework, which outlines essential capabilities for managing UAVs within LAA [icao_utm_framework]. This framework emphasizes the need for global harmonization in airspace management and the integration of UAVs into existing ATM systems. Key capabilities within the ATM system include flight authorization, real-time information sharing, conflict management, and automated traffic de-confliction, all of which must be supported by LAWNs to ensure safe and coordinated UAV operations. By doing so, LAWNs can ensure the safe management of UAVs, complementing the existing aviation infrastructure.
In the United States, the Federal Aviation Administration (FAA) introduced its ATM concept of operations [FAA2022], envisioning a network of UAV service suppliers that provide services such as digital flight planning, airspace management, and UAV tracking. The FAA’s framework operates through cloud-to-cloud data exchanges, eliminating the need for traditional voice communication with air traffic control. A key feature of the FAA system is Remote ID, which mandates UAVs to broadcast unique identification and location data to enable real-time tracking and compliance with airspace restrictions. LAWNs must integrate with these services, ensuring that UAVs are monitored, controlled, and managed in accordance with FAA regulations. To support studies on future parcel delivery by UAV, the Japan Aerospace Exploration Agency (JAEA) has introduced a dedicated simulation platform for the 2030 timeframe, termed the scalable simulator for knowledge of low-altitude environment (SKALE). This framework emulates BVLoS operations above densely populated urban regions. As illustrated in Fig. 10, a representative delivery scenario is configured with a nominal cruise altitude of 120 m. Analysis based on SKALE indicates that current conflict detection and resolution strategies are not yet sufficient for large-scale, highly automated traffic. In particular, the appropriate setting and modelling of separation minima and flight speeds remain open research problems that require further refinement.
Figure 10: The typical UAV delivery workflow of Ref. [oosedo2021unmanned].
Similarly, the Civil Aviation Administration of China (CAAC) has implemented regulations that follow a risk-based approach to managing LAA [caac_low_altitude_service_system_plan_2018]. UAVs are categorized based on their weight and operational risk level, with all UAVs, except the smallest, required to be equipped with unique electronic identification and registered under a real-name system. In addition, LAWNs are required to integrate geofencing capabilities and enforce remote flight restrictions according to the operational limitations set by regulatory authorities. The policies within China are actively driving the LAE development, generating significant demand for high-capacity, scalable LAWNs that can support dense UAV applications in terms of drone delivery, eVTOL air taxis, and remote sensing. To meet these requirements, LAWNs must provide ultra-reliable communication and low-latency connectivity to ensure seamless integration and compliance with regulatory standards in an expanding LAA. Toward this end, the 3rd Generation Partnership Project (3GPP) has released several technical reports (TRs) addressing LAWN communications [3gpp_tr_36777]. Notably, TR 36.777 explores the UAV connectivity by leveraging LTE, highlighting key challenges such as interference and mobility. Furthermore, TR 38.811 and TR 38.821 focus on the integration of 5G new radio (NR) within non-terrestrial networks (NTN), which includes satellites and HAPs.
7.3 Emerging Path Planning Technology
The preceding discussion on low-altitude ATM focuses on macroscopic aspects such as capacity, separation policy, and regulatory coordination. In practice, these system-level concepts are ultimately realized through the individual trajectories flown by each aerial vehicle. Therefore, path planning forms the microscopic counterpart of ATM and directly determines whether a low-altitude airspace can accommodate dense operations without compromising safety or efficiency. A well-designed trajectory must remain compatible with corridor structures, restricted zones, and dynamic flow constraints, while obeying platform limits on energy, payload, and maneuverability. In the context of LAWNs, where numerous heterogeneous UAVs operate in close proximity, deficiencies in path design manifest as conflict hot spots, unnecessary detours, and excessive delay. From a design perspective, most emerging path planning technologies are driven by two primary objectives, namely, trajectory safety and trajectory efficiency [aggarwal2020path]. Specifically, safety-oriented planning emphasizes collision and obstacle avoidance, while compliance with no-fly regions and separation minima. On the other hand, efficiency-oriented planning seeks to reduce mission cost, for example, by shortening flight time and lowering energy consumption through multiobjective criteria. In many recent studies, safety is enforced through hard constraints, whereas efficiency guides the selection of a preferred solution within the feasible set.
For safety-oriented path planning, configuration-based methods remain a central tool, in which the state of a drone, including its position, heading, altitude, and possibly velocity, is embedded into an abstract configuration space in which each point represents a feasible geometric configuration of the vehicle. Physical obstacles, terrain, and no-fly regions are mapped to forbidden subsets of this space, and path planning is formulated as the problem of finding a continuous curve in the free subset that connects the initial and target configurations while obeying kinematic and dynamic constraints. This geometric abstraction allows explicit treatment of vehicle size and maneuverability and provides clear guarantees on collision avoidance once a path in the free configuration space has been obtained. For instance, the work in [67] refined the classical artificial potential field method by introducing a distance-dependent repulsive function, which increases clearance from obstacles and alleviates local minima in dense environments. The authors decomposed the repulsive effect into separate speed and heading components and combined it with a collision prediction mechanism, which yielded smoother trajectories while preserving strong safety margins. The work in [69] employed an improved jump point search model for multirotor UAVs and demonstrated that the resulting paths were both shorter and more maneuverable than those obtained with standard A* search in static urban scenes. To overcome the limitations of two-dimensional grids, the authors in [70] extended jump point search to a three-dimensional representation and proposed modified parent node expansion and pruning strategies, which significantly increased spatial flexibility and computational efficiency, in which the risk, acoustic impact, and logistics cost were incorporated into an A*-based planner for urban delivery missions and showed that trajectories can be shaped to satisfy safety and environmental constraints while retaining practical runtime.
In the context of multiple aircraft, the cooperative and non-cooperative methods can explicitly capture interactions and support online conflict resolution in dynamic environments. The authors in [72] formulated dynamic collision avoidance in high-density urban low-altitude airspace as a Markov decision process and integrated the resulting algorithm into an operational system, achieving real-time conflict management. Moreover, the work in [73] proposed a genetic algorithm-based route replanning method for scenarios with unknown obstacles, where the appearance of new hazards triggers a global search for alternative paths that maintain feasibility and separation. The study in [74] introduced a two-layer reinforcement learning model, in which a local layer performs short-horizon obstacle avoidance and a global layer selects long-horizon routes, enabling simultaneous handling of static and dynamic obstacles with improved path continuity.
By contrast, efficiency-oriented path planning methods build on this safety foundation and place additional emphasis on operating cost and computational complexity. For instance, the work in [77] modeled UAV path planning under partial observability as a partially observable Markov decision process and sought trajectories that minimize cumulative operating cost while accounting for nonlinear dynamics and measurement uncertainty. This formulation provides a principled way to balance performance and robustness when the vehicle has only partial knowledge of its environment. Along this line, the work in [cui2014improved] observed that traditional UAV path planners often ignore mission-specific constraints such as terminal heading and proposed a bidirectional ant-colony-optimization scheme on a grid-based workspace to address this limitation. Ant colonies are initialized near the start and goal configurations and expand simultaneously along predetermined directions, after which a dedicated path-selection mechanism combines and filters candidate routes. Improved pheromone-update rules and successor-selection strategies were introduced to accelerate convergence and reduce the risk of local optima. From a cooperative perspective, the authors in [belkadi2019design] developed a distributed controller for a fleet of UAVs under a multi-agent-systems framework, where online path planning is formulated as a particle-swarm-optimization (PSO) problem solved independently on each vehicle. By minimizing a suitably designed cost function, each local PSO instance determines a trajectory that contributes to formation keeping, target tracking, and collision avoidance, and experiments with fixed and moving targets, external disturbances, and agent loss confirmed that the distributed strategy can maintain coherent fleet behavior and safety without centralized coordination. Overall, emerging path planning technologies complement low-altitude ATM by bridging system-level policies with executable trajectories. Safety-oriented configuration-space methods and cooperative decision frameworks provide reliable guarantees in structured and moderately dynamic airspace, whereas efficiency-oriented approaches enhance scalability and performance in large-scale, high-density LAWNs. The combined use of these approaches is expected to be essential for supporting dense, task-driven operations while maintaining both operational safety and mission effectiveness.
@@ 8.1 Digital Twin-Enabled LAWNs
8.2 Lightweight AI-Enhanced LAWNs
Although integrating large AI models into LAWNs can enable intelligent decision-making, autonomous mission planning, and semantically aware resource allocation on UAVs and edge nodes, practical deployment at the network edge is severely constrained by limited computation, energy budgets, and model size. These constraints make straightforward deployment of current large models infeasible on mobile aerial platforms. Consequently, the design of lightweight and energy-efficient LLM architectures, together with distributed inference frameworks tailored to UAV and edge hardware, remains a central open problem for realizing next-generation intelligent LAWNs.
@@ 8.3 Standardizations for LAWN Deployments
Despite the significant potential of LAWNs in enabling LAE applications, their large-scale adoption is hindered by the lack of standardized frameworks and regulatory guidelines. Key challenges include defining airspace usage protocols, ensuring interoperability among heterogeneous UAV platforms, allocating spectrum for aerial communications, and ensuring compliance with safety and privacy regulations. Standardization efforts must strike a balance between promoting innovation and ensuring regulatory compliance, enabling the reliable, scalable, and secure deployment of LAWNs. Furthermore, the global harmonization of technical, operational, and legal requirements across various regions remains a critical issue that must be addressed to fully unlock the commercial potential of LAWNs.
8.4 3D Interference management
As UAVs operate in 3D airspace, managing interference in LAWNs becomes increasingly complex due to the dynamic nature of both UAV movements and communication channels. Effective 3D interference management is critical to ensure reliable communication, reduce latency, and optimize network capacity. This involves not only minimizing interference between UAVs and ground stations but also managing interactions between UAVs at various altitudes and those operating in adjacent airspaces. Key challenges in 3D interference management include real-time spectrum allocation, adaptive beamforming, and interference mitigation techniques that account for the unique mobility patterns of UAVs.
8.5 Multi-Domain Traffic Management
Multi-domain traffic management (MDTM) seeks to harmonize advanced low-altitude airspace management with conventional air traffic management and emerging space traffic management. LAWNs must share consistent rules, intent information, and safety constraints with vehicles operating across different altitudes and time scales, which is complicated by heterogeneous standards and legacy systems. Therefore, designing unified interfaces and cross-domain conflict resolution mechanisms remains intractable for embedding LAWNs into an integrated air-and-space traffic ecosystem.
8.6 Low-Altitude 4D Trajectory Optimization
Low-altitude four-dimensional (4D) trajectory optimization aims to jointly design three-dimensional flight paths and time profiles so that UAVs satisfy mission deadlines, separation constraints, and airspace capacity limits in a coordinated manner. In dense LAWNs, this problem becomes highly coupled across vehicles and strongly affected by uncertainties in demand, weather, and communication quality. Developing scalable optimization and learning-based methods that can generate conflict-free, energy-efficient 4D trajectories in real time, while remaining compatible with existing ATM and path-planning frameworks, remains a key open research challenge.
8.7 Low-Altitude Traffic Flow Prediction
Low-altitude traffic flow prediction aims to forecast UAV demand and spatiotemporal flow patterns across corridors and sectors, providing critical inputs for capacity planning, congestion mitigation, and conflict management in LAWNs. Compared with traditional air traffic, low-altitude flows exhibit strong nonstationarity, tight coupling with urban activities and weather, and sparse or heterogeneous data sources. Developing reliable data-driven and model-based prediction frameworks that can fuse multi-source information, quantify uncertainty, and operate in real time remains a key open challenge for scalable low-altitude traffic management.
@@ 9 Conclusion
This paper provided a comprehensive survey of LAWNs, focusing on their functional designs and LAA management. We began by comparing LAWNs with traditional aerial communication systems, highlighting their ability to integrate communication, sensing, computation, and control. We then established the foundational principles for LAWN architectures and channel modeling, followed by a review of recent advancements in performance metrics, functional designs, security/privacy preservation, and AI-driven approaches. Finally, we summarized methodologies for LAA structuring, traffic management, and path planning, essential for ensuring scalable and safe LAWN operations.